OpenStudy (anonymous):

Factor completely: 49e^4-49

OpenStudy (anonymous):

I tried setting is up like this: (7e^2)^2 - 7^2, then got (7e^2+7)(7e^2-7) but the answer returned incorrect.

ganeshie8 (ganeshie8):

try this : \[49e^4 - 49 = 49(e^4-1) = 49(e^2+1)(e^2-1) = 49(e^2+1)(e+1)(e-1)\]

OpenStudy (anonymous):

Thanks, that worked. Does x^2-1 always factor out to (x+1)(x-2)? Then is x^2+1 always prime?

OpenStudy (anonymous):

sorry, (x+1)(x-1)

ganeshie8 (ganeshie8):

assume we're living in integers then you're right, x^2+1 is a prime polynomial which cannot be factored

ganeshie8 (ganeshie8):

Notice below : \[x^2 - 1 = x^2-1^2\] use the identity : \[a^2-b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

yw:)